Two gardens are being Fenced in. Both Gardens Will require the same amount of fencing (both Gardens have the same perimeter in meters). One garden is in the shape of a square and the other is in the shape of a equilateral triangle. Each side of the triangle is 9 m longer than each side of the square. Find the dimensions of the garden.
Answer by macston(5194)

s=side of square; s+9m=side of triangle
3(s+9m)=4s
3s+27m=4s Subtract 3s from each side.
27m=s ANSWER: A side of the square measures 27 meters.
s+9m=27m+9m=36m ANSWER 2: A side of the triangle measures 36 meters.
CHECK
Perimeter of square=4(27m)=108m
Perimeter of triangle=3(36m)=108m
perimeter triangle=perimeter square